### v2

# Culture is Not Trivia: Sociocultural Theory for Cultural NLP

Abstract
The field of cultural NLP has recently experienced rapid growth, driven by a pressing need to ensure that language technologies are effective and safe across a pluralistic user base. This work has largely progressed without a shared conception of culture, instead choosing to rely on a wide array of cultural proxies. However, this leads to a number of recurring limitations: coarse national boundaries fail to capture nuanced differences that lay within them, limited coverage restricts datasets to only a subset of usually highly-represented cultures, and a lack of dynamicity results in static cultural benchmarks that do not change as culture evolves. In this position paper, we argue that these methodological limitations are symptomatic of a theoretical gap. We draw on a well-developed theory of culture from sociocultural linguistics to fill this gap by 1) demonstrating in a case study how it can clarify methodological constraints and affordances, 2) offering theoretically-motivated paths forward to achieving cultural competence, and 3) arguing that localization is a more useful framing for the goals of much current work in cultural NLP.
Culture is Not Trivia: Sociocultural Theory for Cultural NLP
Naitian Zhou,1 David Bamman1and Isaac L. Bleaman2 1School of Information, University of California, Berkeley 2Department of Linguistics, University of California, Berkeley {naitian,dbamman,bleaman}@berkeley.edu
1 Introduction
Language and culture are closely linked: language can be conceptualized simultaneously as an artifact of culture as well as a process through which culture is created Ochs (2009). As language technologies become increasingly integrated into the everyday lives of a diverse set of users, it is imperative that they are robust to cultural differences between user bases Hershcovich et al. (2022).
Cultural NLP, sometimes also known as cultural alignment, is a subfield within the NLP and ML communities that has experienced drastic growth in recent years to meet this challenge. Work in cultural NLP usually involves building or evaluating systems that 1) have knowledge of cultural facts and 2) apply this knowledge appropriately in specific situations where cultural knowledge is relevant Adilazuarda et al. (2024); Liu et al. (2024b). This can include building new evaluation benchmarks or fine-tuning datasets that contain cultural knowledge of some kind, either manually Lee et al. (2024); Koto et al. (2024) or automatically from a large corpus Shi et al. (2024); Wang et al. (2024a), or creating systems that generate culturally-relevant output Khanuja et al. (2024).
Most work relies on various proxies for defining both cultural boundaries and cultural objects. Proxies of cultural boundaries commonly include nationality, religion, ethnicity, or other demographic features. Proxies for cultural objects might include culture-specific knowlege of foods, values, or norms Zhou et al. (2024a); Sorensen et al. (2024); Dwivedi et al. (2023). These works constitute an important step forward in understanding how to build fairer, more inclusive language technologies. However, the disparate array of cultural proxies being evaluated is symptomatic of a theoretical gap: to achieve culturally-competent NLP systems, we must make progress towards a clearer, more unified conception of culture, and what it means for the systems we build to be responsive to that.
Fortunately, cultural NLP is not alone in the search for a useful notion of culture, and its theoretical challenges are not new. Dissatisfaction with the coarseness of demographic cultural boundaries led to the second wave of sociolinguistics, which refocused efforts on identifying local cultural meaning within communities of practice Eckert (2012). Larger questions, like the utility of the culture concept, have been debated in fields like sociocultural anthropology, where some researchers have abandoned culture altogether as being essentializing and othering Vann (2013). Indeed, epistemological and empirical tensions as they relate to the study of culture have been grappled with across such fields as anthropology, sociolinguistics, sociology, cultural studies, among many others.
Contributions.
In this paper, we draw on theoretical developments in sociocultural linguistics Bucholtz and Hall (2005) — itself a collection of several adjacent disciplines — to clarify the status of cultural knowledge in building culturally competent NLP systems.
We first review the goals of cultural NLP, and enumerate specific desiderata for culturally-aware language technologies that current works pursue (§2). Then, we highlight recurring difficulties in cultural NLP (§3) by providing a survey of common self-stated limitations in existing papers. We introduce sociocultural linguistics as a field with a useful theoretical framework which we can apply to better understand culture as an object of study (§4), and provide a case study to illustrate how the theory of indexicality can be applied to clarify the distinction between learning cultural knowledge and learning stereotypes (§5). Finally, we discuss the broader implications granted by this understanding of culture, offering two main claims. First, we highlight existing methodological and theoretical gaps in achieving the ambitious goal of cultural competence, and provide theoretically-motivated suggestions for making progress on the task (§6.1). Then, we argue that the goal in cultural NLP might be reasonably understood as localization instead of cultural competence or understanding, providing a more tractable and situated framing with which to build useful NLP systems (§6.2).
2 The goals of cultural NLP
AdaptiveDiscerning
InclusiveNuanced
Figure 1: Two separate spaces of desiderata; the left represents aspects of competence, while the right represents aspects coverage. In each space, there often exists a trade-off between the axes, so that most cultural NLP work falls into the conceptual area that is shaded.
Though the field of cultural NLP does not necessarily agree on a definition of culture, there is general agreement on the goal: to build culturally-competent NLP systems Bhatt and Diaz (2024). Here, we break down this high-level goal into several more specific desiderata that are frequently mentioned in cultural NLP papers. We want our language technologies to be:
Adaptive.
A foundational premise of cultural NLP is that language technologies should be culturally sensitive. In other words, culturally competent language technologies should be responsive to specific cultural contexts when designing their outputs. It would be insufficient for an NLP system to produce the same output for all cultural contexts; many works on bias in NLP have shown and problematized the tendency of language technologies to represent, exaggerate, and perpetuate a hegemonic set of values and structures Voigt et al. (2018); Sheng et al. (2019); Bender et al. (2021).
A wide body of work focuses on assessing whether NLP systems can generate different, appropriate outputs in response to different cultural contexts — by answering value-oriented survey questions in a manner consistent with the target culture, for example ( Cao et al., 2024a; Huang and Yang, 2023). Some works that center adaptation as a value focus on extrinsic evaluation: instead of probing whether language models know specific cultural facts, they test whether NLP systems respond in a way that demonstrates this knowledge Bhatt and Diaz (2024).
Discerning.
At the same time, there is also a desire that NLP systems not perpetuate reductive stereotypes. Past work has demonstrated that users have differing expectations for cultural adaptation Lucy et al. (2024), and that cultural adaptation is not equally desired in all settings. For example, users may want technologies to understand their regional or ethnic dialects, but not generate them Blaschke et al. (2024). This has motivated work on stereotype mitigation ( Jha et al., 2023; Ma et al., 2023), in which datasets of harmful stereotypes are collected in order to evaluate or engineer systems to avoid generating them.
Inclusive.
Cultural NLP values breadth: language technologies should perform well across a large number of cultures. This value is represented by many works in the genre which build benchmarks for a large number of different cultures ( Bhutani et al., 2024); these papers usually use nationality and surveys as tractable ways of achieving large scale ( Zhao et al., 2024; Ramezani and Xu, 2023). Some text mining methods for accumulating cultural knowledge also reflect the emphasis on large-scale, broad coverage Fung et al. (2024); Nguyen et al. (2023).
Nuanced.
In addition to breadth, there is a desire for depth in the form of more granular and extensive cultural understanding. The value of nuance motivates works which build resources for specific languages or contexts Koto et al. (2024); Son et al. (2025); Li et al. (2024b) and explore locally meaningful cultural categories ( Dev et al., 2023). These works might rely on local informants to provide cultural knowledge Koto et al. (2024), trading breadth of coverage for richer cultural knowledge that large-scale survey-based methods cannot capture.
Though these four desiderata are not mutually exclusive, they roughly coalesce into two sets of trade-offs (visualized in fig. 1). The first two desiderata reflect two kinds of cultural competence: the knowledge of how to respond differentially, and the knowledge of when it is appropriate to do so. The second two desiderata reflect an orthogonal value of cultural coverage: we want systems that cover many cultures, as well as systems that cover many aspects of each culture.
3 Recurring troubles
Explicitly stating these desiderata can shed light on the motivations of current work, but they do not themselves offer any answers about what “culture” is. This becomes apparent when we look into the limitations sections of many cultural NLP papers, where we find recurring themes that point to challenges posed by overly narrow definitions of culture. We survey the self-stated limitations of 57 papers from 2022-2024 which explicitly mention culture, as well as the cultural proxies they use. [^1] The full list of papers can be found in the appendix. This is not meant to be an exhaustive survey, but rather illustrative of the general state of the field.
The most commonly cited limitation was one of coverage (40% of papers): the dataset or evaluation being presented was only collected with respect to a small subset of cultures. Partly, this can be explained by the proxies being used for setting cultural boundaries. Of the papers we surveyed, 36% of them used nationality as a demographic proxy Adilazuarda et al. (2024). However, many papers problematize this choice in the limitations section as lacking nuance, since nations are politically defined and not culturally homogeneous Méndez and Naples (2014), and language labels usually reflect a hegemonic notion of a standard variety Lippi-Green (2012).
Another common limitation was a lack of dynamicity (12% of papers): culture is constantly constructed through social negotiation Ochs (2009), but benchmarks are largely static collections of examples or facts Son et al. (2025); Keleg and Magdy (2023); Jin et al. (2024); Li et al. (2024b). In most works, there is no granularity in the temporal dimension, failing to achieve an aspect of the desired nuance. Uncertainty around the definition of culture also limits nuance and inclusivity in cultural technologies, since most papers focus only on a small subset of culturally-relevant objects through proxies like food, etiquette, or values, without a framework to unify them. Roughly 37% of papers directly problematize their choice of a particular cultural proxy as being limited in its ability to represent culture as a whole, or too coarse to capture intragroup variation (28% of papers mention this specifically).
Finally, the tension between adaptation and discernment results in uncertainty about how to address stereotypes in data. Some papers (14%), which are largely intended for use in aligning models, view the potential of collecting stereotypes as cultural knowledge to be a limitation Shi et al. (2024). Other papers explicitly collect stereotypes in order to build systems which can avoid generating them Bhutani et al. (2024).
Other limitations mentioned in various works include an overemphasis on English-language data and methods, the lack of extrinsic evaluation in favor of multiple-choice knowledge tests, the use of pretrained models to construct datasets, and various concerns with crowdsourced or human-annotated data, including the possibility that individual preferences are being construed as cultural ones.
Little progress has been made on rigorously addressing these limitations. In many of these instances, there are questions in clear need of theoretical answers: how do we move past static, global categories when defining culture; how do we conceptualize culture in a way that respects its dynamic and constructed quality; how can we unify different facets of culture; how do we appropriately model and study stereotypes to build fairer systems?
4 A sociocultural solution
The notion of cultural competence is most commonly referenced in social and health services research, where culturally competent care has been encouraged as a way to reduce disparities in care quality and outcomes Alizadeh and Chavan (2016). Similar to cultural NLP, these works face operational challenges in identifying what cultural competence should look like Kirmayer (2012). Often, this literature draws on sociological and anthropological work to resolve these challenges. We will do the same here, focusing our attention on the fields of linguistic anthropology and sociolinguistics, which study language and culture in tandem.
For computational linguists, cultural competence might evoke the notion of linguistic competence, or Gumperz’s more general idea of communicative competence Gumperz (1997): the “knowledge of linguistic and related communicative conventions that speakers must have to initiate and sustain conversational involvement.” Gumperz argues that communication must be understood not only in the context of linguistic systems of grammar, but within a semiotically rich social space. This idea has been widely accepted and refined in both linguistic anthropology and sociolinguistics, [^2] Indeed, Gumperz was greatly influential in the establishment and progress of both these fields. and we take social context as a point of departure for understanding culture.
4.1 Text and context
Insofar as culture can be construed as a structured social phenomenon, it makes sense to understand it as the aspects of (extralinguistic) social context which make themselves interactionally relevant. In Gumperz’s terms, this context contributes to a more general level of sensemaking in an interaction. Edwards (1991) points out that even our linguistic categories are not subject only to cognitive processes, but also to social ones: the semantic category of “bird” might evoke an image of a robin or sparrow in a test-taking setting, but certainly indexes a different one at Thanksgiving dinner. [^3] In the United States, turkey is often a centerpiece in the Thanksgiving meal.
Addressing culture as social context shifts the ambiguity from one term to the other. The question becomes, what do we take to be social context? It is useful to look at the evolution of sociolinguistics as another quantitative discipline in which this question is at the fore. Early sociolinguistic studies focused frequently on sociological categories like socioeconomic class Labov (1966); Guy (2011) or gender Lakoff (1973), placing the speaker as a passive member of an externally-imposed category Eckert (2012). This led to concerns about the limits of coarse macrosociological categories, much like the critique of nationality in cultural NLP today. In response, the second wave of sociolinguistic research incorporated ethnographic methods to better understand local dynamics of language variation, relying on social networks and locally-relevant social categories. While second wave studies focused on local meaning, they still treated social categories as static, an essentializing assumption that equates identity with group affiliation. Third wave studies focus on identity as a performance constructed from a diversity of semiotic resources including, but not limited to, language style. Social context, then, becomes the space within which identity is constructed and performed.
This evolution represents not only theoretical developments in response to empirical challenges in sociolinguistics, but also a steady convergence of ideas with other disciplines. Bucholtz and Hall (2005) provide a well-integrated framework for analyzing language and sociocultural identity in the form of sociocultural linguistics, which synthesizes a convergent set of ideas from across disciplines to analyze language as well as other semiotic practices. Thus, sociocultural linguistics should not be thought of as a single theory of culture, but rather a concordant collection of theories that have seen relative convergence across fields that study language, culture, and society.
We take this framework as the point of departure for the rest of this paper. We describe some foundational concepts (§4.2) and provide a case study for how they can clarify the objects and goals of cultural NLP (§5). Then, we take a broader look at how sociocultural linguistic theory can inform computational work on cultural competence (§6.1) and advance the goals of cultural NLP (§6.2).
4.2 A primer on sociocultural linguistics
Bucholtz and Hall (2005) lay out five core principles of sociocultural linguistics. The terminology they use centers on the concept of identity, defined as “the social positioning of self and other.” If we take the analogy of identity as a social position, we might consider culture to be the broader landscape within which identities are located. Through this lens, it becomes clear that individual identities not only reflect, but also constitute culture, giving shape to the cultural terrain. Thus, sociocultural linguistics provides a useful way of conceptualizing the cultural system more generally. [^4] While Bucholtz and Hall (2005) is nominally centered around identity and interaction, they in fact discuss culture more generally at various points (e.g., when introducing partialness).
Emergence.
Identity emerges through interaction. This is the view that language does not come from culture, but rather that culture is constituted through linguistic (and other forms of) interaction. This draws on, among others, the ideas of identity performance Butler (1988) and audience design Bell (1984). Emergence supports a more nuanced representation of culture as one that is dynamic, and a more complex notion of adaptation that supports the idea that even an individual can inhabit multiple cultural roles. This offers one solution to the challenge of defining cultural categories: it may make sense to instead induce cultural categories latent in the data.
Positionality.
Identity includes multiple levels of categories, including macro-level demographics, locally-specific distinctions, and contextually-specific stances and styles. This is a more inclusive and nuanced notion of culture that speaks to one of the core limitations of current work. National identity is only one level at which identity occurs, and the idea of positionality insists that we understand more granular categories of identity as well, including ones that are local to a specific community or even an interaction.
Indexicality.
Identity is constructed through an indexical process involving signs and their conceptual referents Silverstein (2003); Eckert (2008). Indexicality offers a mechanism through which culture is constructed; it is the process of drawing links between linguistic (and other) forms and social meaning. These can play into cultural ideologies about language, construct stances local to specific interactions, consist of overt references to identity, and more; this provides a unified mechanism through which we can conceptualize culture. In section 5, we explore an example of how indexicality provides a useful theoretical account of stereotype in cultural NLP.
Relationality.
Identity takes on social meaning in relation to other identities. This provides a useful way of conceptualizing culture that aligns with findings in machine learning: contrastive learning of feature spaces often result in stronger representations than supervised learning among predefined categories. Similarity and difference are not the only relations available within this framework, which also includes authentication-denaturalization and authorization-illegitimization, among others. If the cultural space is structured through these relations, computational methods might benefit from considering how to encode them.
Partialness.
Finally, any account of culture is necessarily incomplete, since it is itself situated contextually in relation to the subject it describes. Indeed, a person’s identity at a given point in time may be partially deliberate, partially habitual (and subconscious), partially attributable to perception, partially conditioned by the interactional context, and partially subject to the ideologies that surround the interaction. That there is no single ground truth is a troubling statement for those who want to build robust, generalizable systems. However, it also provides a certain freedom from a positivist mirage. Instead, researchers and system designers are encouraged to think more critically about their position, and the assumptions encoded in the technology they build, with respect to the users whom these systems impact.
5 Case study: culture from text
In §4.1, we distinguish between language (the text) as the traditional object of study in linguistics and various ways of assessing culture as the surrounding social context. It has become essentially paradigmatic within NLP that we should expect to derive extratextual information (such as world models, for example) from training on text alone. It is reasonable, then, that there is a vein of cultural NLP work that mines for facts about specific identities from culturally-centered discourse on social media or the Internet Rao et al. (2025); Fung et al. (2024); Nguyen et al. (2023). In some cases, large language models are prompted to generate specific culturally-relevant scenarios Qiu et al. (2024). In many of these papers, authors note the dangerous potential for extracting biased or stereotyping information. How should we understand the epistemic status of these surfaced facts as cultural knowledge? Labov (1972, 314) defines stereotype as the linguistic forms which are subject to metapragmatic discussion: that is, the signs whose meanings are actively discussed. In this section, we apply indexical theory to demonstrate that these works, in fact, can only learn stereotypes.
Papers that mine for cultural facts aim to construct an indexical field from unstructured web text. These systems take as input online documents or discussions about cultural differences (e.g., a guide to dining etiquette in India; Rao et al., 2025) and extract indexical mappings resembling this form:
In cultural group, belief is widely accepted.
This is effectively a mapping between the space of beliefs and the cultural groups that they index. Indexicality can occur at different levels of social awareness; a first-order index evidences membership in a group. For example, the use of “pop” over “soda” might index membership in the population of Midwestern U.S. English speakers. [^5] In the Midwest, it is widely accepted that fizzy, sugary drinks are called “pop.” However, higher-order indices occur as these associations themselves become embedded in cultural ideology. As such, the understanding that Midwesterners say “pop” is in and of itself a piece of cultural knowledge Eckert (2008). [^6] In the U.S., it is widely accepted that Midwestern U.S. English speakers use “pop” over “soda.”
5.1 Stereotypes all the way down
The implication of this idea is that works which study cultural discourse are primarily studying the stereotypes embedded in the ideologies of the groups that generated this data, and only incidentally studying the cultures that are the objects of discourse. The status of these cultural facts as stereotypes does not depend on whether the group generating the discourse is the same as the group serving as the subject: the fact that these are cultural associations being discussed rather than observed classifies them as stereotypes. In fact, groups can, and often do, have both positive and negative stereotypes for themselves Leavitt et al. (2015); Coffman (2014); Sinclair et al. (2006); Pickett et al. (2002).
By applying the theory of indexical order, we gain clarity on the aspects of culture being studied. In the case of papers that mine cultural knowledge from cultural discourse, we find the subjects of study to be different from what we initially assumed. We are not learning about a diverse set of international cultures, but rather the world-view of the text authors, situated in a specific interactional context (perhaps posting about culture shock).
This is not just a matter of naming, nor a dismissal of the utility of these datasets. Instead, indexical theory clarifies the extent to which they are useful. It shows that these datasets exclude, by construction, cultural knowledge that is not subject to metapragmatic discussion. It shows that higher-order indices can still be useful because their meanings are tied to the lower-order ones from which they arise. But it also illustrates complications that we must contend with: higher-order indices might persist even when lower-order ones are no longer as salient. “Authentic” Pittsburghers, for example, might be described as unpretentious, hospitable, sports-loving, etc. But this style originally indexed the immigrant-heritage, working-class history of the formerly industrial city, an identity that may not necessarily apply to its current residents, many of whom work in the health-care or higher-education sectors Johnstone (2014).
This is also not to say that first-order indices are inaccessible through computational methods. For example, computational sociolinguistic work successfully identifies first-order social meaning by using platform metadata like Twitter geolocation Grieve et al. (2019) or subreddit Zhang et al. (2017); Lucy and Bamman (2021) to measure sociolinguistic variation in linguistic features (often lexical or semantic) and the sociocultural meaning they index (like locale or community).
5.2 Indexical values are contextual
It is also a mistake to assume that a given style from a given speaker always indexes the same thing, because the indexical value is also dependent on context, and interactionally interpreted.
Chun (2007), for example, provides an account of a “foreign speaker” language style as deployed by Asian American high schoolers. She notes how this style can be employed both as accommodation to foreign speakers (e.g., a child speaking to her immigrant parents) and as mockery (e.g., between two peers at school). Sometimes, quotatively, an utterance can even fulfill both roles depending on the interactional frame through which it is interpreted. The social meaning of an utterance is determined situationally within a specific interactional context.
6 Paths forward
Sociocultural linguistics paints a picture of culture as a complex, dynamic system through which sense-making occurs. It is one that has proven useful in accounting for and describing how semiotic systems are constructed and deployed for social action in everyday interactions.
6.1 Culturally competent NLP
But there exists a gap between this model of culture and our current computational methods for approaching culture. There is opportunity for NLP work to fill in these gaps.
Sociocultural linguistic theory tells us that culture is emergent, and cultural NLP acknowledges that culture is a dynamic process, but currently our datasets are limited to static snapshots of cultural artifacts. It may be fruitful to instead analyze discursive sequences in which cultural knowledge is suggested or contested. When and how are norms enforced in interaction? How is cultural knowledge shared, and how is it taken up by the rest of the community? As an example, consider this interaction between two Latina high school students from Mendoza-Denton (2008): {adjustwidth}-0.5cm
Lupe: ¿Qué me ves?
(What are you looking at?)
Patricia: Tschhh, don’t EVEN talk to me in Spanish, ‘cause your Spanish ain’t all that.
Through contextual information like the participants’ posture, make-up, and social networks (including the fact that they are rival gang members), we can understand the setting of this interaction: Patricia has interpreted Lupe’s question to be a claim to authenticity. But through the interaction itself we can see the cultural process in action: as Mendoza-Denton (2008) notes, Lupe asserts her Mexican-ness symbolically through her use of Spanish. Through both her assertion and Patricia’s contestation, the social importance of Spanish is reinforced as indexing their Mexican identities. Analogous computational work might study comment threads for these kinds of interactions, and additionally incorporate contextual mechanisms like flairs or voting that users can employ to express affiliation or pass judgment on platforms like Reddit Gaudette et al. (2021).
Sociocultural theory tells us that culture is positional, operating at multiple levels of identity and often composed of features from many different styles, but current methods impose coarse, usually unidimensional, boundaries like nationality on cultural categories. Relationality and indexicality offer mechanisms through which cultural sensemaking occurs — how can we better model positionality as a contextually legible field of identities by identifying instances of cultural categories being constructed in relation to other categories, or identities being assembled by combining different indexical signs? Castelle (2022) suggests that modern language models can be usefully conceptualized more generally as effective learners of semiotic systems; how might we build systems that learn representation spaces for other kinds of meaning beyond semantics, like social or discursive meaning?
Indexicality also motivates the need for datasets that are contextually rich: culture is the combination and construction of different semiotic resources that make reference to social meaning, yet our methods are deployed on datasets that largely consist of decontextualized text. Data that contains social context in other forms (e.g., metadata or other kinds of world state) could be one way of addressing this limitation Nguyen (2025). For example, the STAC corpus Asher et al. (2016) consists of dialogue situated in a game scenario, and includes information about the game state and actions. This places linguistic interaction within a broader context; future works might extend this paradigm to other, more socially relevant metadata.
Indexical fields also exist beyond text, reaching into other modalities in the form of gesture, prosody, and even extralinguistic semiotic systems like fashion Chun (2007) and make-up Mendoza-Denton (2008). Not only is it important to represent non-text modalities to capture culture, but combining modalities can also be a promising direction to learning social meaning Zhou et al. (2024b).
There is also much theoretical work at hand to account for how a software system might differ from a human in how it is taken up as an interlocutor in interaction. Creating systems that perfectly replicate human behavior is neither desirable nor felicitous. Consider this podcast transcript introducing the findings of a scientific paper:
{adjustwidth}
-0.5cm-0.5cm
Host A: Think about those old Hollywood films, the ones your grandma might watch.
Host A: Do those performances feel different than what you might see in movies today?
Host B: Hm, yeah I guess they do. It’s, like, more dramatic. The emotions are way more, out there?
This serves the discursive purpose of simultaneously motivating a finding and establishing rapport with the listener by drawing on the presenter’s personal experience. However, if this same script is generated with an LLM, [^7] As, indeed, it was, by NotebookLM Google (2024). the social action becomes infelicitous. The LLM has no grandmother, whose past movie-going experiences are being imagined and described. Instead, the audience must reinterpet this sequence as a post-hoc rationalization of the source material that is about to be presented, failing to motivate the finding or establish rapport. Not all semiotic resources available to humans are available to the language technologies we build.
6.2 Localized NLP
All told, we are far from building culturally competent systems, given these clear and pressing theoretical and methodological gaps. But cultural NLP also faces more immediate goals, which are perhaps more central to the field as it is currently configured. We want to create, e.g., web agents that will not make food purchases that violate religious dietary laws Qiu et al. (2025) or image generation models that show the local currency when displaying money Khanuja et al. (2024). These goals are mediated by constraints on resources: user studies and qualitative interviews might provide richer cultural data, but are more expensive to scale. Do we need to achieve cultural competence in the general sense for these more immediate applications? And how do we navigate trade-offs between desiderata?
Many would argue that machine translation systems have not yet achieved linguistic competence (and this is perhaps an easier case to make in the multilingual setting). Yet, individual software applications have been internationalized long before MT achieved even its most recent success. When building systems that accommodate more users, a more useful, immediate framing might be one of localization rather than cultural competence. Understanding the task at hand as building localized NLP applications helps us locate ourselves in the space of desiderata (fig. 1).
Localization is tractable because it forces us to specify the application domain, constraining the relevant depth of knowledge. Localized translations are generated only for the necessary text within an application; culturally localized systems can focus on the domain-specific nuances of cultural knowledge. Sociocultural approaches to culture are contextual and situated, and localization forces us to evaluate cultural performance in a situated application setting.
Localization also forces us to enumerate our audience, constraining and making explicit the coverage of our systems. Localized translations are not provided for an arbitrary, unconstrained set of languages or an arbitrary set of text. Furthermore, a website that offers its interface in, e.g., “Spanish” rarely allows users to choose a specific regional dialect, even though different varieties of Spanish often show lexical and syntactic variation. This is a pragmatic choice, but also an ideological one about which language varieties to support, and it is better that the ideological choices be made explicitly and transparently. In the cultural setting, this can also make the choice of cultural boundaries less arbitrary. If the goal is to build a culturally localized healthcare chatbot, for example, differing levels of medical literacy may be a more salient cultural boundary with more actionable interventions than something like nationality.
Finally, localization forces us to consider the NLP system as an interlocutor in the human-computer interaction. While many existing cultural knowledge benchmarks probe large language models removed from the specific context of how they will be used, approaching the task as localization forces us to define the expected behavior within a given application context. Developers of a recipe application might improve user experience by offering culture-specific ingredient substitutions She et al. (2024), but a healthcare application might benefit from adopting a stance of cultural humility instead of potentially stereotyping or stigmatizing adaptation Lekas et al. (2020). Defining the bounds of expected cultural performance specifies where we want a particular application to lie in the discerning / adaptive space.
Thus, localization allows us to focus on particulars that may be more tractably implemented and evaluated in real-world systems today. Principles from user-centered Lowdermilk (2013) and participatory design Caselli et al. (2021), particularly as applied to machine learning contexts Ayobi et al. (2023); Tan et al. (2024), provide footholds for NLP practitioners who wish to identify the cultural features which are relevant to a given application context.
7 Conclusion
It is important to build language technologies that are responsive to cultural values. However, the current field of cultural NLP has not found agreement on what it means to model culture, settling instead for a wide array of cultural proxies for both categories of identity and categories of indexical features. In this paper, we deconstruct the goals of cultural NLP and highlight how recurring discomforts in current work are illustrative of a lack of theoretical alignment. We propose drawing on convergent theoretical insights from a variety of social-scientific disciplines which have centered the study of culture in the context of language and other semiotic systems.
When studying such a complex, multifaceted, and dynamic object as culture, it is equally challenging and imperative that the object of study be well-defined. We demonstrate how sociocultural linguistics provides a useful theoretical framework that treats culture as an enacted process, not a static artifact. We explore the implications of this: we show that learning cultural facts through metapragmatic discourse is limited to learning about stereotypes; we make the case that building culturally competent computational systems requires a dynamic model of culture as a process, not a collection of trivia, and that sociocultural linguistics provides a powerful model of this process, but methodological and theoretical gaps still loom large; finally, we argue that many current works in cultural NLP can be usefully reframed as localization, which encourages situated, participatory design and evaluation of systems.
The growth of cultural NLP reflects the more general state of natural language processing. Until recently, NLP has been largely preoccupied with learning the semantic meaning of textual symbols. But other fields of linguistics have long established that the world around us cannot be extricated from the words we produce and interpret. Gumperz argued fifty years ago that communicative competence reaches beyond grammatical knowledge. As computational methods have become more powerful in representing textual semantic meaning, it becomes both tractable and necessary to consider other kinds of meaning, like sociocultural meaning, as equally important objects of study.
8 Limitations
Though we highlight certain representative works in making the case for a theory-led approach to cultural NLP, this paper is not meant to be a survey of the field. Adilazuarda et al. (2024) and Liu et al. (2024b) offer good overviews of recent work.
We also introduce the specific theoretical framework of sociocultural linguistics. Though other theories of culture certainly exist, we focus on sociocultural linguistics for its linguistically-oriented approach to culture (see §4 for detailed discussion about this).
Acknowledgments
This work was supported by funding from the National Science Foundation (Graduate Research Fellowship DGE-2146752).
We thank Laura Sterponi, Kent Chang, Dan Hickey, Nikita Mehandru, Lucy Li, and many others for their thoughtful discussion and feedback.
Appendix A List of surveyed works
Paper Limitations
Good Night at 4 pm?! Time Expressions in Different Cultures Shwartz (2022) proxy, individual, multilingual
Probing Pre-Trained Language Models for Cross-Cultural Differences in Values Arora et al. (2023) proxy, survey
GD-COMET: A Geo-Diverse Commonsense Inference Model Bhatia and Shwartz (2023) intrinsic, stereotype
Assessing Cross-Cultural Alignment between ChatGPT and Human Societies: An Empirical Study Cao et al. (2023) language as culture, proxy
Sociocultural Norm Similarities and Differences via Situational Alignment and Explainable Textual Entailment CH-Wang et al. (2023) coarseness, individual, coverage
Toward Cultural Bias Evaluation Datasets: The Case of Bengali Gender, Religious, and National Identity Das et al. (2023) coverage
Building Socio-culturally Inclusive Stereotype Resources with Community Engagement Dev et al. (2023) context, coverage, multilingual
NORMSAGE: Multi-Lingual Multi-Cultural Norm Discovery from Conversations On-the-Fly Fung et al. (2023) stereotype, culture is dynamic
Multilingual Language Models are not Multicultural: A Case Study in Emotion Havaldar et al. (2023) coverage, coarseness
Culturally Aware Natural Language Inference Huang and Yang (2023) coarseness, stereotype
SeeGULL: A Stereotype Benchmark with Broad Geo-Cultural Coverage Leveraging Generative Models Jha et al. (2023) context, coarseness, subjectivity
Multi-lingual and Multi-cultural Figurative Language Understanding Kabra et al. (2023) coarseness, unmarked culture
DLAMA: A Framework for Curating Culturally Diverse Facts for Probing the Knowledge of Pretrained Language Models Keleg and Magdy (2023) proxy, coarseness, crowdsourced
Evaluating the Diversity, Equity, and Inclusion of NLP Technology: A Case Study for Indian Languages Khanuja et al. (2023) proxy, language as culture, coarseness
NormMark: A Weakly Supervised Markov Model for Socio-cultural Norm Discovery Moghimifar et al. (2023) coverage
FORK: A Bite-Sized Test Set for Probing Culinary Cultural Biases in Commonsense Reasoning Models Palta and Rudinger (2023) proxy, coverage, annotator, reductive
Knowledge of cultural moral norms in large language models Ramezani and Xu (2023) coverage, culture is dynamic
NLPositionality: Characterizing Design Biases of Datasets and Models Santy et al. (2023) proxy
Geographical Erasure in Language Generation Schwöbel et al. (2023) multilingual
Modeling Cross-Cultural Pragmatic Inference with Codenames Duet Shaikh et al. (2023) stereotype
Cross-Cultural Analysis of Human Values, Morals, and Biases in Folk Tales Wu et al. (2023) coverage, multilingual
Cross-Cultural Transfer Learning for Chinese Offensive Language Detection Zhou et al. (2023a) multilingual, coarseness
Cultural Compass: Predicting Transfer Learning Success in Offensive Language Detection with Cultural Features Zhou et al. (2023b) proxy, coverage, coarseness
NormBank: A Knowledge Bank of Situational Social Norms Ziems et al. (2023a) coverage, stereotype
Multi-VALUE: A Framework for Cross-Dialectal English NLP Ziems et al. (2023b) proxy
Investigating Cultural Alignment of Large Language Models AlKhamissi et al. (2024) coverage, proxy, reductive, coarseness
Extrinsic Evaluation of Cultural Competence in Large Language Models Bhatt and Diaz (2024) proxy, multilingual
Bridging Cultural Nuances in Dialogue Agents through Cultural Value Surveys Cao et al. (2024a) language as culture, pluralism
Cultural Adaptation of Recipes Cao et al. (2024b) proxy, coarseness, coverage
The Echoes of Multilinguality: Tracing Cultural Value Shifts during LM Fine-tuning Choenni et al. (2024) language as culture, proxy, coarseness
Massively Multi-Cultural Knowledge Acquisition & LM Benchmarking Fung et al. (2024) LLM-derived
ViSAGe: A Global-Scale Analysis of Visual Stereotypes in Text-to-Image Generation Jha et al. (2024) subjectivity, annotator, reductive, proxy, discerning
KoBBQ: Korean Bias Benchmark for Question Answering Jin et al. (2024) subjectivity, proxy
CLIcK: A Benchmark Dataset of Cultural and Linguistic Intelligence in Korean Kim et al. (2024) proxy, coarseness, coverage, stereotype, discerning
The PRISM Alignment Dataset: What Participatory, Representative and Individualised Human Feedback Reveals About the Subjective and Multicultural Alignment of Large Language Models Kirk et al. (2024) discerning, preference v morality, exploited annotators
IndoCulture: Exploring Geographically-Influenced Cultural Commonsense Reasoning Across Eleven Indonesian Provinces Koto et al. (2024) culture is dynamic, coverage
KorNAT: LLM Alignment Benchmark for Korean Social Values and Common Knowledge Lee et al. (2024) culture is dynamic, unmarked culture, annotator, intrinsic
CultureLLM: Incorporating Cultural Differences into Large Language Models Li et al. (2024a) intrinsic
CULTURE-GEN: Revealing Global Cultural Perception in Language Models through Natural Language Prompting Li et al. (2024c) intrinsic, coarseness
Are Multilingual LLMs Culturally-Diverse Reasoners? An Investigation into Multicultural Proverbs and Sayings Liu et al. (2024a) proxy, coverage
Cultural Alignment in Large Language Models: An Explanatory Analysis Based on Hofstede’s Cultural Dimensions Masoud et al. (2025) culture is dynamic, multilingual, proxy
Having Beer after Prayer? Measuring Cultural Bias in Large Language Models Naous et al. (2024) coarseness, coverage
Cultural Commonsense Knowledge for Intercultural Dialogues Nguyen et al. (2024) crowdsourced, LLM-derived, stereotype
Evaluating Cultural and Social Awareness of LLM Web Agents Qiu et al. (2025) multilingual, coverage, context
NormAd: A Framework for Measuring the Cultural Adaptability of Large Language Models Rao et al. (2025) proxy, coarseness, culture is dynamic, multilingual, coverage
DOSA: A Dataset of Social Artifacts from Different Indian Geographical Subcultures Seth et al. (2024) coverage, intersectionality, multilingual
CultureBank: An Online Community-Driven Knowledge Base Towards Culturally Aware Language Technologies Shi et al. (2024) multilingual, unmarked culture, stereotype
KMMLU: Measuring Massive Multitask Language Understanding in Korean Son et al. (2025) copyright, intrinsic
Value Kaleidoscope: Engaging AI with Pluralistic Human Values, Rights, and Duties Sorensen et al. (2024) LLM-derived
Navigating Cultural Chasms: Exploring and Unlocking the Cultural POV of Text-To-Image Models Ventura et al. (2025) coverage, based on existing model
CRAFT: Extracting and Tuning Cultural Instructions from the Wild Wang et al. (2024a) multilingual
SeaEval for Multilingual Foundation Models: From Cross-Lingual Alignment to Cultural Reasoning Wang et al. (2024b) coverage, intrinsic
Not All Countries Celebrate Thanksgiving: On the Cultural Dominance in Large Language Models Wang et al. (2024c) survey, coverage
Benchmarking Machine Translation with Cultural Awareness Yao et al. (2024) proxy
RENOVI: A Benchmark Towards Remediating Norm Violations in Socio-Cultural Conversations Zhan et al. (2024) multilingual
WorldValuesBench: A Large-Scale Benchmark Dataset for Multi-Cultural Value Awareness of Language Models Zhao et al. (2024) culture is dynamic, proxy, individual
Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge Zhou et al. (2024a) coverage, multilingual, crowdsourced
Table complete
Appendix B Category key
Category Description
coverage Makes note of insufficient or limited coverage.
proxy Notes the limitations of the chosen proxy.
coarseness Mentions that cultural boundaries are too coarse.
multilingual Mentions that only one (or a limited number) of languages are studied.
stereotype Mentions concerns about learning or perpetuating stereotypes.
culture is dynamic Mentions limitations in capturing the dynamicity of culture.
intrinsic Mentions the limitations of only performing intrinsic evaluation.
language as culture Mentions the limitations of treating language as cultural boundaries.
crowdsourced Mentions concerns about crowdsourced data.
discerning Mentions that not all cultural attributes should be treated the same way.
LLM-derived Mentions concerns that some part of the dataset was generated by LLMs.
reductive Mentions concerns that culture is reduced to the proxies chosen.
annotator Mentions concerns about human-annotated data.
unmarked culture Mentions concerns with “universal” cultural attributes or values.
individual Mentions that individual values in the data may not be reflective of cultural ones.
subjectivity Mentions that cultural annotations are subjective.
context Notes that the system or datasets lacks social or situational context.
survey Notes the limitations of relying on survey data for cultural knowledge.
pluralism Mentions concerns with choosing which value to align to in pluralistic situations.
preference v morality Notes that aligning to preferences may not be desirable behavior.
exploited annotators Notes that annotators providing preference data do not generally share in the benefits.
intersectionality Mentions limitations in covering intersectional identities.
copyright Notes that some data points were removed due to copyright.
based on existing model Notes that automated evaluation is based on an existing model.
Table complete